OpenStudy (anonymous):

i(sq root of -98) - (sq root of 98)

OpenStudy (anonymous):

first off \[\sqrt{98}=\sqrt{49\times 2}=7\sqrt{2}\] so work with that one

OpenStudy (anonymous):

\[i(7\sqrt{2}i-7\sqrt{2})=7\sqrt{2}i^2-7\sqrt{2}i=-7\sqrt{2}-7\sqrt{2}i\]\]

OpenStudy (anonymous):

i assumed you mean the whole thing times i. if it is just the first term then my answer is wrong

OpenStudy (anonymous):

if it is just the first term multiplied by i the answer is \[-14\sqrt{2}\]